OpenStudy (anonymous):

Write the equation of the circle with center (-4, 11) and containing the point (5, -1). Use the ^ key for the exponents. Write your answer as the example: (x+2)^2+(y-5)^2=49

OpenStudy (anonymous):

(x+4)^2 + (y-11)^2 = 169

OpenStudy (anonymous):

Mmm...don't agree Assuming point (5,-1) lies ON the circle, radius R: (x+4)^2 + (y-11)^2 = R^2 Sub in x=5 and y=-1 giving (5+4)^2 + (-1-11)^2 = R^2, or 9^2 + (-12)^2 = 81 + 144 = 225 = R^2 So R =sqrt 225 = 15 Equation of circle therefore (x+4)^2 + (y-11)^2 = 225